How much money should be deposited today in an account that earns 5% compounded semiannually so that it will accumulate to $8000 in 11 years?
Answer by rfer(10432)

FV=P(1+r/n)^n*t
P=FV/(1+r/n)^n*t
P=8000/(1.025)^22
P=8000/1.721571398
P=$4647